C = 100 + 0.8Yd
investment (I) = 200
govt. exp. (G) = 200
export = 1000
import function(M) = 20+0.2Y

(a) Calculate equilibrium income Y
(b) Calculate value of multiplier and equilibrium income if import function (M) = 20+0.3Y
1
Expert's answer
2020-03-30T07:20:23-0400

"a) Y=100+0.8Yd+200+200+1000-(20+0.2Y)"


"Y=1480+0.6Y"

"Y-0.6Y=1480"

"0.4Y=1480"

"Y=3700"

"b) Y=100+0.8Yd+200+200+1000-(20+0.3Y)"

"Y=1480+0.5Y"

"0.5Y=1480"

"Y=2960"